A body thrown up with a finite speed is caughback after 4 sec. The speed of the body withwhich it is thrown up is

Grade:11

1 Answers

sree lakshmi
3 years ago
total time =4sec
time required to reach max height =4/2=2sec
velocity at max height v =0
initial velocity u=?
acceleration = -g
then v=u+at 
          =u-gt=0
         u=gt 
            =9.8×2=19.6m/s
 or approx. = 20m/s